OpenStudy (anonymous):

Given the function f(x) = −2x + 4, find x if f(x) = 20.

OpenStudy (anonymous):

-8=x

OpenStudy (anonymous):

Need me to show you how I did it?

jimthompson5910 (jim_thompson5910):

f(x) = -2x+4 20 = -2x+4 ... replace f(x) with 20 Now solve for x

OpenStudy (anonymous):

Or that.

OpenStudy (anonymous):

` yes please help me , thanks

jimthompson5910 (jim_thompson5910):

f(x) = -2x+4 20 = -2x+4 20-4 = -2x+4-4 16 = -2x 16/(-2) = (-2x)/(-2) -8 = x x = -8